a small mass is suspended from a long thread so as to form a simple pendulum. the period T of oscillation depends on the m the length L of the thread and the acceleration g of free fall at the place concerned T= KM^x L^y G^z where x,y,z are unknown and K =211 (1) Find the values of X Y Z (2) Hence derive an expression relating T M G and L
1
Expert's answer
2019-11-07T09:10:28-0500

"T= km^x l^y g^z"

"[T]= [m]^x [l]^y [g]^z"

"T= [M]^x [L]^y [LT^{-2}]^z"

"-2z=1\to z=-\frac{1}{2}"

"y-z=0\to y=\frac{1}{2}, x=0"

Thus,


"T=2\pi\sqrt{\frac{l}{g}}"